OpenStudy (hockeychick23):

If the mean is subtracted from every number in a data set, and the differences are used to form a new data set, that data set must have a mean: A. equal to the mean of the original data set B. equal to zero C. smaller than the mean of the original data set D. larger than the mean of the original data set E. cannot be determined by the information given

OpenStudy (hockeychick23):

@ganeshie8

OpenStudy (campbell_st):

well just try it.... pick 3 numbers say 5, 8, 17 whats the mean..? what is the mean of the differences..?

OpenStudy (campbell_st):

try another set and see if your conclusions hold true

OpenStudy (hockeychick23):

I tried that and i got a smaller number (c) but i wasn't sure if thats always true

OpenStudy (hockeychick23):

I tried that and i got a smaller number (c) but i wasn't sure if thats always true

OpenStudy (campbell_st):

(5 + 8 + 17)/3 = 10 5 - 10 = -5 8 - 10 = -2 17 - 10 = 7 add them up and they add to zero... so go and test it with another data set

OpenStudy (hockeychick23):

(1+4+9+10)/4= 6 1-6= -5 4-6= -2 9-6= 3 10-6= 4 ok i got zero for these

OpenStudy (hockeychick23):

so would that make it B? @campbell_st

OpenStudy (campbell_st):

so it seems to be a thing... so the sum of the differences is 0 now divide by the number of terms... to get the mean in my example 0/3 = 0 your example 0/4 = 0

OpenStudy (hockeychick23):

so am i right then?